find the value of the missing term(s) of each of the following proportions.
A. 5/3 = N/12
B. 7/2 = 42/N
C. N/2 = 50/N

Answer by cleomenius(959)

1. 5/3 = n/12 n = 20
2. 7/2 = 42/n n = 10
3. n/2 = 50/n n*2 = 100 ; n = 10
Cleomenius